OpenStudy (anonymous):

solve \(3x^2+3xy-5y^2=55\)

OpenStudy (anonymous):

x,y are real numbers

OpenStudy (anonymous):

This a curve. What do you mean by solve it.

OpenStudy (kainui):

Print it out, put the paper in water or some other kind of "solvent" to dissolve it. Problem "solved"

OpenStudy (anonymous):

It is a hyperbola

OpenStudy (anonymous):

i believe we are loocing for pairs \(x,y\) i have found such pairs for integers but i am concerned with none integers,real numbers

OpenStudy (anonymous):

^^^looking

OpenStudy (anonymous):

irs should satisfy the hyperbola

OpenStudy (anonymous):

There are infinite numbers of (x,y) that satisfy your equation. They are lie on the hyperbola drwan by wolframalpha http://www.wolframalpha.com/input/?i=plot3d+3x^2%2B3xy-5y^2%3D55

OpenStudy (primeralph):

@Jonask You should review multivariable functions.

OpenStudy (anonymous):

for exam purposes i want to get that only with a calculator,how do i proceed

OpenStudy (anonymous):

ohh this is a number theory class,not calculus

OpenStudy (primeralph):

You have x and y, so there'll be many solutions for each.

OpenStudy (anonymous):

@primeralph

OpenStudy (primeralph):

Yeah, either way you need to tell us what you're solving.

OpenStudy (anonymous):

The points (4, 1) (-5, 1) satisfy the equation

OpenStudy (anonymous):

well basically all solutions for example>>>\(\dfrac{27+\sqrt{69}}{2}\)

OpenStudy (anonymous):

i have found integer solutions here http://math.stackexchange.com/questions/739752/integer-solutions-of-the-equation-of-the-form-ax2bxycy2-m#

OpenStudy (anonymous):

concerned with none integer solutions,ie real numbers

OpenStudy (anonymous):

NONE INTEGER SOLUTIONS

OpenStudy (anonymous):

more generally the book gives solutions \(\pm(\dfrac{17\pm\sqrt{29}}{2})(2+\dfrac{1\pm\sqrt{29}}{2})^{2n} \cup (\dfrac{41\pm7\sqrt{29}}{2})(2+\dfrac{1\pm\sqrt{29}}{2})^{2n}\) \[n\]is integer

OpenStudy (anonymous):

Here is a list of all integer solutions betwee -100 and 100 {{-100, 53}, {-53, 28}, {-47, -53}, {-25, -28}, {-25, 13}, {-12, -13}, {-9, 4}, {-5, -4}, {-5, 1}, {-4, -1}, {4, 1}, {5, -1}, {5, 4}, {9, -4}, {12, 13}, {25, -13}, {25, 28}, {47, 53}, {53, -28}, {100, -53}}

OpenStudy (anonymous):

Here is a list of all integer solutions between -1000 and 1000 {{-620, -701}, {-620, 329}, {-291, -329}, {-213, 113}, {-100, -113}, {-100, 53}, {-53, 28}, {-47, -53}, {-25, -28}, {-25, 13}, {-12, -13}, {-9, 4}, {-5, -4}, {-5, 1}, {-4, -1}, {4, 1}, {5, -1}, {5, 4}, {9, -4}, {12, 13}, {25, -13}, {25, 28}, {47, 53}, {53, -28}, {100, -53}, {100, 113}, {213, -113}, {291, 329}, {620, -329}, {620, 701}}

OpenStudy (anonymous):

@eliassaab how can we generalise the integer solutions,and how did you find other integer solutions?

OpenStudy (anonymous):

I wrote a Mathematica program to generate the list. Change p and the program below and execute Clear[f, x, y] p = 1000; f[x_, y_] = 3 x^2 + 3 x y - 5 y^2 - 55; Q = {}; For[ m = -p, m <= p, m++, For[ n = -p, n <= p, n++, If[f[m, n] == 0, Q = Append[Q, {m, n}]]]] Q

OpenStudy (anonymous):

ok,i understand integer solutions,sir do you know how we can find---real number solutions,i mean non integers

OpenStudy (anonymous):

Take a any real number, replace y by a. You will have a quadratic in x, solve it in terms of a and you are done if your discriminant is positive or zero

OpenStudy (anonymous):

There are infinite number of them.

OpenStudy (anonymous):

That is what you get \[ \left\{\left\{x\to \frac{1}{6} \left(-\sqrt{3} \sqrt{23 a^2+220}-3 a\right)\right\},\left\{x\to \frac{1}{6} \left(\sqrt{3} \sqrt{23 a^2+220}-3 a\right)\right\}\right\} \]

OpenStudy (anonymous):

yes \[\Delta=69\] but what steps are necessary to do this \(O_{69}=[1,\omega_{69}]=[1,\frac{1+\sqrt{69}}{2}]\)

OpenStudy (anonymous):

You can see that the discriminant of the quadratic is always positive

OpenStudy (anonymous):

so you take a as a constant,then treat it as a normal quadratic of the form \[\alpha x^2+\beta x+c=0\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i would be interested i finding \(a\) as well

OpenStudy (anonymous):

This a list of some non integer solutions \[ \left( \begin{array}{cc} \frac{1}{6} \left(\frac{3}{13}-\frac{3 \sqrt{12401}}{13}\right) & -\frac{1}{13} \\ \frac{1}{6} \left(-\frac{9}{130}-\frac{\sqrt{11154621}}{130}\right) & \frac{3}{130} \\ \frac{1}{6} \left(-\frac{24}{65}-\frac{6 \sqrt{77581}}{65}\right) & \frac{8}{65} \\ \frac{1}{6} \left(-\frac{87}{130}-\frac{3 \sqrt{1245781}}{130}\right) & \frac{29}{130} \\ \frac{1}{6} \left(-\frac{63}{65}-\frac{\sqrt{2818929}}{65}\right) & \frac{21}{65} \\ \frac{1}{6} \left(-\frac{33}{26}-\frac{3 \sqrt{50501}}{26}\right) & \frac{11}{26} \\ \frac{1}{6} \left(-\frac{102}{65}-\frac{42 \sqrt{1626}}{65}\right) & \frac{34}{65} \\ \frac{1}{6} \left(-\frac{243}{130}-\frac{\sqrt{11606709}}{130}\right) & \frac{81}{130} \\ \frac{1}{6} \left(-\frac{141}{65}-\frac{3 \sqrt{326769}}{65}\right) & \frac{47}{65} \\ \frac{1}{6} \left(-\frac{321}{130}-\frac{3 \sqrt{1327109}}{130}\right) & \frac{107}{130} \\ \frac{1}{6} \left(-\frac{36}{13}-\frac{2 \sqrt{30369}}{13}\right) & \frac{12}{13} \\ \end{array} \right) \]

OpenStudy (anonymous):

They are all of the form \[ \left\{\frac{1}{6} \left(-\sqrt{3} \sqrt{23 a^2+220}-3 a\right),a\right\} \] Vary a and you get the other cooerdinates

OpenStudy (anonymous):

If you want of find a as well, put x = b and solve the quadratic in terms of b

OpenStudy (anonymous):

\[ \left\{\left\{y\to \frac{1}{10} \left(3 b-\sqrt{69 b^2-1100}\right)\right\},\left\{y\to \frac{1}{10} \left(\sqrt{69 b^2-1100}+3 b\right)\right\}\right\} \] Here you have to chose b to have the discriminant positive

OpenStudy (anonymous):

hmmm ok, this is what i wanted ,so logically this would mean we are looking for values where the level curve's achieve a value of 55 for \(f(x,y)=3x^2+3xy-5y^2\),thank you @eliassaab

OpenStudy (anonymous):

by the way you said \(a,b\) are real numbers,it is therefore hard to generate such numbers,that will gve a positive dirsciminant

OpenStudy (anonymous):

if it was integers then it would be easy to generate such integers

OpenStudy (anonymous):

psh, ur all mathletes